OpenStudy (anonymous):

find the domain of the following rational expression -7xy^2/28x^5y

OpenStudy (anonymous):

the domain is all x values for which the expression is defined, so what you need to ask is: "at what values for x is the expression not defined?"

OpenStudy (anonymous):

in the case of rational expressions, the points where it will be undefined are where you attempt to divide by zero, so for what x value(s) does the denominator equal zero?

OpenStudy (anonymous):

would the answer be = y^2 5y

OpenStudy (anonymous):

to find where the expression is undefined set 28x^5y=0 if you think about it, the x is the only factor that matters because x^(any power) will be a number not equal to zero and so the expression will be defined there, so what value of x makes 28x^5y=0?

OpenStudy (anonymous):

\[x=x25y\]

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i am not rying to sound like a dummy but algebra is one of my weak subjects im suffering

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

you do not need to worry about the y in this case because no value of y will make the denominator zero so what value of x will make the denominator zero? you want a number not another expression

OpenStudy (anonymous):

4

OpenStudy (anonymous):

think about the zero property

OpenStudy (anonymous):

is (-7xy^2)/(28x^(5y)) your whole expression?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so ignore the ^5y what would you have to multiply 28 by to get zero?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

exactly, so if you substitute zero in for x you get \[\frac{ -7(0)y^2 }{ 28(0)^5y }\] which means the numerator and denominator are both zero when x is zero. when this happens it creates a "hole" in the graph of the expression or function which means it is discontinuous and not defined there, in this case the expression is defined for all x except 0

OpenStudy (anonymous):

so what x=0

OpenStudy (anonymous):

i meant to say so the domain is x=0

OpenStudy (anonymous):

the domain would be \[{ R, x \neq 0 }\]

OpenStudy (anonymous):

all real numbers except zero

OpenStudy (anonymous):

so the answer is x -0

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

the domain is the set of all x values for which an expression is defined, this expression is not defined at x = 0 so the answer is \[x \neq 0\] and not x = 0

OpenStudy (anonymous):

what is the name of that sign youre using

OpenStudy (anonymous):

not equal to